Hello, I could really use your help with this question.
What is the expression for the height of a cone in terms of its radius and volume?
Thanks, Amanda

Answer by khwang(438)

I think you mean a right circular cone.
From the formula of the volume, (A is the area of the base)
V = 1/3 A h = 1/3 pi r^2 * h

so, h = 3V/(pi r^2)
That's it.
Kenny